# Temporal Sepsis Modeling: a Fully Interpretable Relational Way

### v1

@@ IV-C Univariate variable importance
The analysis of the discretization allows us to understand why this aggregate is informative. Small values of $V_{1}$ produce intervals relatively pure of Sepsis=1 while the intervals between 26 and 36 are relatively pure of Sepsis=0. The reading of $P(X|C)$ per interval, and for each variable, is informative to understand the result of the univariate classifications.
IV-D Local variable importance
To analyze local behavior (for each example), it is possible to compute the Shapley values of all the input variables of the trained classifier using the method described in [19] and available in the Khiops library V11. This allows to have the variable importance example by example (local importance). We illustrate this information (XAI) in the Table III.
@@ IV-D Local variable importance
Table III: Illustration of the knowledge provided by the local importance
IV-E Global variable importance
The global importance of the variables is given in column 3 of the Table II. These importance are defined for each variable as the average on all the train samples of the absolute values of the local Shapley values described in the previous section.
IV-F Counterfactuals
@@
A counterfactual explanation might be “If your income had been \$10000 higher, then your credit would have been accepted” [37, 21, 20]. A semi-factual is a special-case of the counterfactual in that it conveys possibilities that “counteract” what actually happened, even if the outcome does not change [1]: “Even if your income had been \$5000 higher your credit would still be denied” (but closer to being accepted).
Within the framework of counterfactual reasoning, we here used the method described in [22] and the notion of trajectory. Indeed, the study of counterfactual trajectories a posteriori is of great interest, as it also makes it possible to identify when a trajectory is approaching the frontier or pass trough the frontier.
In this paper, the trajectory of a counterfactual is the minimum of changes in the input vector to change the predicted class to the opposite [^2] A notebook that performs this process for the Khiops library is given in the Github for reproducibility.
In the case of Sepsis, one could be interested, for example, when examining Patients :
@@
V Note on state-of-the-art in Machine Learning for Early Sepsis Prediction
Recent advances in machine learning have led to several commercial systems for early prediction of sepsis, evaluated in heterogeneous clinical settings. Among the most representative are InSight® [7], Sepsis ImmunoScore™ [34], NAVOY® Sepsis [29] and VFusion™ Sepsis [36] report high performance, with AUC values generally between 0.83 and 0.91. However, as highlighted by the PREVIA [35] retrospective study and in line with existing meta-analyses [10], direct comparison of these systems remains challenging due to substantial differences in patient populations, prediction horizons, variables used and validation protocols.
Note on reproducibility and validity
The initial dataset (MIMIC-III) which contains 40000 patients as well as the 3940 patients used in the experiments are available on the GitHub (https://anonymous.4open.science/r/sepsis-project-5DBD/README.md). The code to train the classifiers (Table I), output the local importance (Table III), the counterfactuals, etc, is also available in the same Github. In this way, all the results are fully reproducible. Moreover one author of this paper is a doctor specialized on Sepsis who validated the found results.
VI Conclusion
In this study, we presented an innovative approach for the early detection of sepsis by exploiting electronic medical data in relational form. By combining the relational representation of patient trajectories, the propositionalization technique based on the MDL method, and a selective Bayesian classifier, we succeeded in obtaining a model that is both effective and intrinsically interpretable. This approach not only improves prediction accuracy but also provides a clear understanding of the key factors contributing to sepsis detection, thereby facilitating acceptance by clinicians.
The experimental results highlight the relevance of this approach, particularly its ability to produce understandable features and adapt to different prediction horizons. In addition, the flexibility of relational representation opens up prospects for the integration of multimodal data and the management of patient trajectories of varying sizes.
@@ VI Conclusion
In the context of sepsis, counterfactuals offer the possibility of identifying the changes needed in clinical variables to shift a prediction from positive to negative (or vice versa). For example, by adjusting certain physiological parameters of a patient, it becomes possible to understand which factors most influence the model’s decision. This approach promotes a better understanding of the underlying mechanisms and can guide medical interventions by proposing alternative trajectories to improve care.
This study, which incorporates counterfactual reasoning and local explanation techniques, helps to increase transparency and confidence in predictive models in clinical settings. Continuing this work could contribute to more personalized, accurate, and interpretable medicine, particularly in the fight against complex syndromes such as sepsis.

### v2

@@ 4.3 Univariate variable importance
4.4 Local variable importance
To analyze local behavior (for each example), it is possible to compute the Shapley values of all the input variables of the trained classifier using the method described in [18] and available in the Khiops library V11. This allows to have the variable importance example by example (local importance). We illustrate this information (XAI) in the Table 2.
@@ 4.4 Local variable importance
Table 2: Illustration of the knowledge provided by the local importance
4.5 Global variable importance
The global importance of the variables is given in column 3 of the Table 3. These importance are defined for each variable as the average on all the train samples of the absolute values of the local Shapley values described in the previous section.
4.6 Counterfactuals
@@
A counterfactual explanation might be “If your income had been \$10000 higher, then your credit would have been accepted” [19].A semi-factual is a special-case of the counterfactual in that it conveys possibilities that “counteract” what actually happened, even if the outcome does not change [1]: “Even if your income had been \$5000 higher, your credit would still be denied” (but closer to being accepted). Within the framework of counterfactual reasoning, we here used the method described in [20] and the notion of trajectory. Indeed, the study of counterfactual trajectories a posteriori is of great interest, as it also makes it possible to identify when a trajectory is approaching the frontier or pass trough the frontier. In this paper, the trajectory of a counterfactual is the minimum of changes in the input vector to change the predicted class to the opposite [^2] A notebook that performs this process for the Khiops library is given in the Github for reproducibility. In the case of Sepsis, one could be interested, for example, when examining Patients :
@@
Note on classifier performance and on reproducibility
Recent advances in machine learning have led to several commercial systems for early prediction of sepsis, evaluated in heterogeneous clinical settings. Among the most representative are InSight® [6], Sepsis ImmunoScore™ [32], NAVOY® Sepsis [27] and VFusion™ Sepsis [34] report high performance, with AUC values generally between 0.83 and 0.91. Highlighted by the PREVIA [33] retrospective study and in line with existing meta-analyses [9], direct comparison of these systems remains challenging due to substantial differences in patient populations, prediction horizons, variables used, and validation protocols. However, we give, on GitHub (due to space considerations), a supplementary material related to this paper: a comparison of the results obtained versus 4 other models (Random Forest, CatBoost, XGboost and a LSTM).
The initial dataset (MIMIC-III), which contains 40000 patients, as well as the 3940 patients used in the experiments, is available on GitHub. The code to train the classifiers (Table 3), output the local importance (Table 2), the counterfactuals, etc, is also available in the same GitHub. In this way, all the results are fully reproducible. Moreover, one of the authors of this paper is a doctor specializing in Sepsis who validated the findings.
5 Conclusion
In this study, we presented an innovative approach for the early detection of sepsis by exploiting electronic medical data in relational form. By combining the relational representation of patient trajectories, the propositionalization technique based on the MDL method, and a selective Bayesian classifier, we succeeded in obtaining a model that is both effective and intrinsically interpretable. This approach not only improves prediction accuracy but also provides a clear understanding of the key factors contributing to sepsis detection, thereby facilitating acceptance by clinicians. The experimental results highlight the relevance of this approach, particularly its ability to produce understandable features and adapt to different prediction horizons. In addition, the flexibility of relational representation opens up prospects for the integration of multimodal data and the management of patient trajectories of varying sizes.
@@ 5 Conclusion
In the context of sepsis, counterfactuals offer the possibility of identifying the changes needed in clinical variables to shift a prediction from positive to negative (or vice versa). For example, by adjusting certain physiological parameters of a patient, it becomes possible to understand which factors most influence the model’s decision. This approach promotes a better understanding of the underlying mechanisms and can guide medical interventions by proposing alternative trajectories to improve care. This study, which incorporates counterfactual reasoning and local explanation techniques, helps to increase transparency and confidence in predictive models in clinical settings. Continuing this work could contribute to more personalized, accurate, and interpretable medicine, particularly in the fight against complex syndromes such as sepsis.
Appendix
Khiops
Q R S=#Var Acc Train Auc train Acc Test Auc Test Rob AUC Size
10 5 5 0.7835 0.8647 0.7716 0.8444 0.98 1
100 86 44 0.9332 0.9813 0.9132 0.9737 0.99 1
1000 872 98 0.9530 0.9913 0.9360 0.9833 0.99 1
10000 8849 198 0.9703 0.9962 0.9337 0.9819 0.99 1
100000 72331 272 0.9916 0.9991 0.9406 0.9784 0.98 1
Random Forest
Q R S=#Var Acc Train Auc train Acc Test Auc Test Rob AUC Size
10 - 9 1.0 1.0 0.7831 0.8281 0.83 1.80
100 - 99 1.0 1.0 0.9155 0.9757 0.98 2.25
1000 - 996 1.0 1.0 0.9383 0.9817 0.98 10.16
10000 - 8964 0.9819 0.9985 0.9360 0.9785 0.98 45.27
100000 - 38743 0.9357 0.9841 0.9086 0.9725 0.99 142,43
CatBoost
Q R S=#Var Acc Train Auc train Acc Test Auc Test Rob AUC Size
10 - 9 0.8906 0.9575 0.7899 0.8592 0.90 1.80
100 - 99 0.9964 0.9999 0.9269 0.9802 0.98 2.25
1000 - 860 0.9992 0.9999 0.9406 0.9850 0.98 8.77
10000 - 2662 1.0 1.0 0.9406 0.9833 0.98 13.44
100000 - - - - - - - -
XgBoost
Q R S=#Var Acc Train Auc train Acc Test Auc Test Rob AUC Size
10 - 9 0.9934 0.9997 0.7899 0.8527 0.85 1.80
100 - 88 1.0 1.0 0.9246 0.9786 0.98 2.0
1000 - 607 1.0 1.0 0.9452 0.9847 0.98 6.19
10000 - 1159 1.0 1.0 0.9360 0.9834 0.98 5.85
100000 - 1176 1.0 1.0 0.9337 0.9832 0.98 4.32
LSTSM
- - 34 0.913 0.968 0.868 0.939 0.97 H=256
- - 34 0.908 0.967 0.877 0.936 0.97 H=128
- - 34 0.920 0.972 0.874 0.945 0.97 H=64
- - 34 0.913 0.967 0.877 0.932 0.96 H=32
- - 34 0.894 0.953 0.852 0.920 0.97 H=16
Table 3: Results versus the number of Q variables to be constructed and classifiers (Train 90%, Test 10%.)
Rob Auc: Robustness on the AUC ($\mbox{AUC}_{\mbox{Test}}$ / $\mbox{AUC}_{\mbox{Train}})$)
Size: #Var used by ”RF or CB or XGB” for a the same value of $Q$ / #Var used by khiops for a given value of $Q$
CatBoost did not succeed in dealing with the 100000 Variables on the computer used (HP EliteBook 655 15.6 inch G10 Notebook PC, AMD Ryzen 5 PRO 7530U with Radeon Graphics, 2000 MHz, 6 cores(s), 12 processeur(s) logique(s), RAM 32Go)
LSTM: One hidden layer, H: Number of neurons, Optimiser: Adam, Learning Rate: 0.001, Number of epochs=1000